OpenStudy (anonymous):

CAN ANYONE HELP ME???? Which expressions could complete this equation so that it has one solution? 4(4x-3)-6x=______ Choose exactly two answers that are correct. A. 12x-12-2x B. 9x-12 C. 4x+2(3x-6) D. 4(2x+3)-24

OpenStudy (anonymous):

@green_1 could you help me with this if no one else does?

Directrix (directrix):

4(4x-3)-6x is not an equation. It is an algebraic expression. Would you check the problem in your text to see if the problem is correctly posted? Thanks.

OpenStudy (anonymous):

i missed the = and _____ sorry didn't notice till you said something... oops

OpenStudy (anonymous):

a

Directrix (directrix):

@heyhina808 Please show the work for your direct answer. Thanks.

OpenStudy (anonymous):

it also says i need two answers not just one @heyhina808

Directrix (directrix):

And, option 'a' is not one of them. It leads to an identity.

Directrix (directrix):

Test B: 4(4x-3)-6x= 9x-12 16x - 12 -6x = 9x -12 10x - 12 = 9x -12 x = 0 Option B has a single solution of 0. There is still one more to find.

MsBrains (ms-brains):

C is one of them, I'm pretty sure..

Directrix (directrix):

Let's try option D 4(4x-3)-6x= 4(2x+3)-24 16x - 12 - 6x = 8x + 12 -24 10x - 12 = 8x - 12 10x = 8x 2x = 0 x = 0 0 is a single solution.

MsBrains (ms-brains):

I think the answers are A and C, but I could be wrong. I'm not completely sure @raventayler13

Directrix (directrix):

>C is one of them, I'm pretty sure.. Let's check: 4(4x-3)-6x= 4x+2(3x-6) 16x - 12 - 6x = 4x + 6x - 12 10x - 12 = 10x - 12 This is an identity and has infinitely many solutions. It is not an answer to the posted question.

Directrix (directrix):

@Ms-Brains Try B and D.

Directrix (directrix):

>>I think the answers are A and C, but I could be wrong. @Ms-Brains Let's see some work to back up the claim.

MsBrains (ms-brains):

I did it all in my head, And they both come up to 10x-12. But you know what, I'm probably wrong. @Raventayler13 He's probably right. :)